# Learning Noise-Robust Stable Koopman Operator for Control with Hankel DMD

### v5

@@ III Numerical Experiments
    Cartpole System – Evaluated on the task of stabilizing the pole in the upright position.
@@ III-A Autonomous Van der Pol Oscillator
where $\mu$ is set to 1.0. The Van der Pol Oscillator is a nonlinear system characterized by a non-conservative force that can either dampen or amplify oscillations depending on the state. Since our framework relies on the discrete-time dynamical system, the continuous-time version of the van der Pol system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.1. The training dataset is generated by running simulations for 100 steps, also with a time step of $\Delta t$ of 0.1. After simulation, Gaussian noise is added to the states, with 10 logarithmically spaced noise levels from $10^{-3}$ to $10^{-1}$. For models trained on a single trajectory, the initial condition is sampled from a normal distribution with a mean of $[-1.0,-1.0]$ and a standard deviation of $[0.05,0.05]$. For models trained on multiple trajectories, 50 initial conditions are uniformly sampled from the range $[-1,1]$. To ensure robustness and generalizability, we created 10 sets of training data using different random seeds. Figure 2 illustrates one set of training trajectories and test trajectories within the $x_{1},x_{2}$ plane.
@@
We compare the proposed framework with the following versions of DMD:
@@
Figure 3a compares the reconstruction performance of various methods based on dynamic mode decomposition (DMD) when trained on a single trajectory of the Van der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The dissipative Koopman model with polyflow basis functions (Dissipative-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating a superior precision. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure 3b presents the mean squared error (MSE) trajectory over 100 time steps for the Van der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean error for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The dissipative Koopman model with polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the dissipative Koopman model but performs better than some of the other methods. The figures demonstrate that the dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
@@
Figure 4 compares the prediction errors for multiple trajectories of the Van der Pol (VDP) system across different noise levels. Panel (a) shows the training error, while panel (b) shows the test error. The dissipative Koopman model with polyflow basis functions consistently shows lower errors across various noise levels, demonstrating its robustness and accuracy in handling multiple trajectories.
@@
Lifting Function Identity Identity Polyflow RBF Identity Polyflow RBF Polyflow RBF Polyflow
@@
Mean Normalized Training Error 0.0010 0.1198 0.1211 0.0697 0.1079 0.1210 NaN NaN 0.0547 0.0898 0.0417
@@
Mean Normalized Test Error 0.0010 0.1315 0.1326 0.1028 0.1268 0.1325 NaN NaN 0.0796 0.1085 0.0590
@@ III-B CartPole System
Let us consider the CartPole System,
@@
The CartPole system consists of a cart that moves along a track, balancing a pole attached to it via a pivot. The dynamics of the system are described by the equations above, where $x_{1}$ represents the position of the cart, $x_{2}$ represents the angle of the pole, $x_{3}$ represents the velocity of the cart, $x_{4}$ represents the angular velocity of the pole, and $u$ is the control input force applied to the cart. The parameters of the system include the mass of the cart ($m_{c}$), the mass of the pole ($m_{p}$), the length of the pole ($l$), and the acceleration due to gravity ($g$). For our framework, the continuous-time CartPole system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.05. The training dataset is generated by running simulations for 200 steps and contains 50 different trajectories. After the simulation, Gaussian noise is added to the states with noise level of 0.1. The initial conditions for the CartPole system are randomly sampled with the position ($x_{1}$) from $[-1,1]$, the angle ($x_{2}$) from $[-\pi/2,\pi/2]$, the velocity ($x_{3}$) from $[-0.1,0.1]$, and the angular velocity ($x_{4}$) from $[-0.1,0.1]$. Control inputs are generated using exponentially decaying sine waves of randomly sampled frequencies, amplitudes and phase angles.
@@ Results
Figure 5 compares the Model Predictive Control (MPC) costs for the CartPole system using three different models under a noise level of 0.100. The models compared are EDMD with RBF basis functions (eDMD-RBF), the Dissipative Koopman model with Polyflow basis functions (Dis-PF), and the Standard Koopman model with Polyflow basis functions (Std-PF). The violin plots illustrate the distribution of MPC costs across various initial states. The results indicate that both the Dissipative Koopman model with Polyflow basis functions (Dis-PF) and the Standard Koopman model with Polyflow basis functions (Std-PF) achieve relatively similar MPC costs. Both models show stable control performance across different initial states, demonstrating robustness in control. In contrast, the EDMD with RBF basis functions (eDMD-RBF) exhibits higher variability in MPC costs and generally incurs higher average MPC costs compared to the Dissipative and Standard Koopman models. This suggests that while the EDMD-RBF model provides some level of stability, it does so at the expense of higher control costs and less consistency. Overall, the comparison reveals that both the Dissipative and Standard Koopman models with Polyflow basis functions offer superior and more stable control performance compared to the EDMD-RBF model under the noise level of 0.1. This highlights the effectiveness of the Polyflow basis functions in achieving robust control in MPC applications.
@@ III-C Mountain Car
Mountain Car Continuous Version 0 is an environment from the Gymnasium library [47] where a car is tasked with reaching the top of a mountain. The control action is a force applied along the x-axis with bounds $[-1.0,2.0]$. The observed state consists of the car’s position along the x-axis and its velocity.
We apply our learning framework to this environment under the assumption that the governing equation is unknown, and we have access only to noisy measurements of state-action sequences. For data generation, we use uniformly distributed random actions within the action space bounds and uniformly distributed initial positions between $[-0.6,-0.4]$ with zero initial velocity. The episode terminates upon reaching the target or after 1000 steps, with each step representing 0.02s. We collect 200 episodes of varying step lengths.
To preprocess the dataset, we normalize the states and actions by subtracting their mean values and dividing by their standard deviations. Additionally, we introduce Gaussian noise with mean 00 and variances $\sigma^{2}=\{0.01,0.1,0.5\}$ to create three different training sets. A separate test dataset is generated following the same procedure.
@@
Figure 6: Test trajectory prediction errors for the Mountain Car, Lunar Lander, and Panda environments under unknown dynamics. For these experiments, a Polyflow order of 2 was used for the lifted state construction, and all models were trained with training data subject to 0.1 noise levels. Optimization was performed using the SOAP optimizer [53].
@@ III-D Lunar Lander
The Lunar Lander environment is a Gymnasium environment [47] in which an agent controls a lander’s thrusters to achieve a soft landing on a designated pad. The state space originally comprises the lander’s position, velocity, angle, and angular velocity, while the action space consists of continuous thrust controls.
@@ III-E Panda Emika Franka Robot
Figure 8: Effectiveness of dissipative roll-out under noise. Heatmaps show predictive test performance across different noise levels and sequence lengths.
@@ IV Iterative Data Augmentation
Figure 9 visualizes the trajectory of the CartPole system in three different scenarios: no data augmentation, one-time data augmentation, and twice data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. Performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (one-time augmentation) and 67.8 (twice augmentation).
@@
Figure 10 presents violin plots that illustrate the distribution of Model Predictive Control (MPC) costs for the CartPole system on two different tasks, seen and unseen, and iterative models under a noise level of 0.1. The “seen” task refers to scenarios where input bounds are set between $[-20,20]$, while the “unseen” task refers to tighter input bounds of $[-10,10]$. Data augmentation was performed using closed-loop trajectories from the seen task. It is important to note that iterative classical learning [56] cannot be easily generalized to a different task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative learning. The results indicate that both seen and unseen control tasks benefit significantly from iterative learning, with successive iterations leading to lower and more stable MPC costs. This shows the effectiveness of iterative learning in improving the overall performance of the system.
@@ V Ablation Studies
  • Stabilized formulation of Koopman operator in Section II-G,
@@
V-A Dissipative Koopman Model versus Standard Koopman Model
Figure 13a compares the performance of the dissipative Koopman model and the standard Koopman model in different noise levels. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure 11 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the standard Koopman matrix move outside the unit circle, leading to a loss of stability, while the dissipative Koopman model never loses stability during training.
V-B Learning Rate Scheduler
Figure 13b compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels. For low noise, the cyclic learning rate scheduling achieves a 20% reduction in training error and a 16.67% reduction in test error, with a 9.09% reduction in time taken compared to the constant learning rate. In medium noise scenarios, the cyclic learning rate scheduling shows a 16.67% improvement in training error, 14.29% in test error, and an 8.70% improvement in time taken. However, at high noise levels, cyclic learning rate scheduling underperforms compared to the constant learning rate, with higher training and test errors, as well as slightly longer training times. These results indicate that while the cyclic learning rate scheduler is beneficial at low and medium noise levels, it is less effective in high-noise environments.
V-C Optimizer Switching
Figure 13c examines the effect of switching optimizers on the model’s performance. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching from the Adam optimizer to LBFGS after 4000 epochs in a 5000-epoch training cycle resulted in about a 25% error reduction compared to the Adam-only optimizer at low and medium noise levels, demonstrating the efficacy of LBFGS during fine-tuning. However, this advantage diminished at higher noise levels, where the differences between the optimizers became less pronounced. In experiments where optimizers were switched periodically every 500 epochs, results varied, but overall, periodic switching helped minimize the error in certain cases, though it did not consistently outperform the Adam-LBFGS approach. Notably, using only LBFGS led to significantly higher errors, suggesting that while LBFGS is beneficial for fine-tuning, it tends to settle in local minima and does not further improve the error.
V-D Progressive roll-out length versus Constant roll-out length
Figure 13d compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. Progressive roll-out scheduling shows a 20% improvement in mean training errors and a 16.67% improvement in mean test errors compared to the constant roll-out length, indicating better learning and generalization. Although the time taken for training with progressive roll-out scheduling is 6.94% lower, the overall training duration remains slightly higher due to the incremental increase in roll-out length. These results highlight the effectiveness of progressive roll-out scheduling in enhancing model performance.

### v6

@@ III Numerical Experiments
    Cartpole System – Evaluated on the task of stabilizing the pole in the upright position from [52]
@@ III-A Autonomous Van der Pol Oscillator
where $\mu$ is set to 1.0. The Van der Pol Oscillator is a nonlinear system characterized by a non-conservative force that can either dampen or amplify oscillations depending on the state. Since our framework relies on the discrete-time dynamical system, the continuous-time version of the van der Pol system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.1. The training dataset is generated by running simulations for 100 steps, also with a time step of $\Delta t$ of 0.1. After simulation, Gaussian noise is added to the states, with 10 logarithmically spaced noise levels from $10^{-3}$ to $10^{-1}$. For models trained on a single trajectory, the initial condition is sampled from a normal distribution with a mean of [1.0,1.0][-1.0,-1.0] and a standard deviation of $[0.05,0.05]$. For models trained on multiple trajectories, 50 initial conditions are uniformly sampled from the range [1,1][-1,1]. To ensure robustness and generalizability, we created 10 sets of training data using different random seeds. Figure fig. 2 illustrates one set of training trajectories and test trajectories within the $x_{1},x_{2}$ plane.
@@
We compare the proposed framework with the following noise-robust variants of DMD:
@@
Figure fig. 3a compares the reconstruction performance of various methods based on dynamic mode decomposition (DMD) when trained on a single trajectory of the Van der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The dissipative Koopman model with polyflow basis functions (Dissipative-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating a superior precision. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure fig. 3b presents the mean squared error (MSE) trajectory over 100 time steps for the Van der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean error for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The dissipative Koopman model with polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the dissipative Koopman model but performs better than some of the other methods. The figures and table I demonstrate that the dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
@@
Figure fig. 4 compares the prediction errors for multiple trajectories of the Van der Pol (VDP) system across different noise levels. Panel (a) shows the training error, while panel (b) shows the test error. The dissipative Koopman model with polyflow basis functions consistently shows lower errors across various noise levels, demonstrating its robustness and accuracy in handling multiple trajectories.
@@
Lifting
Function
Identity Identity Polyflow RBF Identity Polyflow RBF Polyflow RBF Polyflow
@@
Mean Normalized
Training Error
0.0010 0.1198 0.1211 0.0697 0.1079 0.1210 NaN NaN 0.0547 0.0898 0.0417
@@
Mean Normalized
Test Error
0.0010 0.1315 0.1326 0.1028 0.1268 0.1325 NaN NaN 0.0796 0.1085 0.0590
@@ III-B CartPole System
Let us consider the CartPole System [52],
@@
The CartPole system consists of a cart that moves along a track, balancing a pole attached to it via a pivot. The dynamics of the system are described by the equations above, where $x_{1}$ represents the position of the cart, $x_{2}$ represents the angle of the pole, $x_{3}$ represents the velocity of the cart, $x_{4}$ represents the angular velocity of the pole, and $u$ is the control input force applied to the cart. The parameters of the system include the mass of the cart ($m_{c}=4kg$), the mass of the pole ($m_{p}=1kg$), the length of the pole ($l=1m$), and the acceleration due to gravity ($g=9.81m/s^{2}$). For our framework, the continuous-time CartPole system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.05. The training dataset is generated by running simulations for 200 steps and contains 50 different trajectories. After the simulation, Gaussian noise is added to the states with noise level of 0.1. The initial conditions for the CartPole system are randomly sampled with the position ($x_{1}$) from [1,1][-1,1], the angle ($x_{2}$) from $\pi/2,\pi/2]$, the velocity ($x_{3}$) from [0.1,0.1][-0.1,0.1], and the angular velocity ($x_{4}$) from [0.1,0.1][-0.1,0.1]. Control inputs are generated using exponentially decaying sine waves of randomly sampled frequencies, amplitudes and phase angles.
@@ Results
Figure fig. 5 compares the Model Predictive Control (MPC) costs for the CartPole system using three different models under a noise level of 0.100. The models compared are EDMD with RBF basis functions (eDMD-RBF), the Dissipative Koopman model with Polyflow basis functions (Dis-PF), and the Standard Koopman model with Polyflow basis functions (Std-PF). The violin plots illustrate the distribution of MPC costs across various initial states. The results indicate that both the Dissipative Koopman model with Polyflow basis functions (Dis-PF) and the Standard Koopman model with Polyflow basis functions (Std-PF) achieve relatively similar MPC costs. Both models show stable control performance across different initial states, demonstrating robustness in control. In contrast, the EDMD with RBF basis functions (eDMD-RBF) exhibits higher variability in MPC costs and generally incurs higher average MPC costs compared to the Dissipative and Standard Koopman models. This suggests that while the EDMD-RBF model provides some level of stability, it does so at the expense of higher control costs and less consistency. Overall, the comparison reveals that both the Dissipative and Standard Koopman models with Polyflow basis functions offer superior and more stable control performance compared to the EDMD-RBF model under the noise level of 0.1. This highlights the effectiveness of the Polyflow basis functions in achieving robust control in MPC applications.
@@ III-C Mountain Car
Mountain Car Continuous Version 0 is an environment from the Gymnasium library [53] where a car is tasked with reaching the top of a mountain. The control action is a force applied along the x-axis with bounds [1.0,2.0][-1.0,2.0]. The observed state consists of the car’s position along the x-axis and its velocity.
We apply our learning framework to this environment under the assumption that the governing equation is unknown, and we have access only to noisy measurements of state-action sequences. For data generation, we use uniformly distributed random actions within the action space bounds and uniformly distributed initial positions between [0.6,0.4][-0.6,-0.4] with zero initial velocity. The episode terminates upon reaching the target or after 1000 steps, with each step representing 0.02s. We collect 200 episodes of varying step lengths.
To preprocess the dataset, we normalize the states and actions by subtracting their mean values and dividing by their standard deviations. Additionally, we introduce Gaussian noise with mean 0 and variances $\sigma^{2}=\{0.01,0.1,0.5\}$ to create three different training sets. A separate test dataset is generated following the same procedure.
Test prediction errors are shown in fig. 6(a), MPC trajectories in fig. 7(a), and noise sensitivity across sequence lengths in fig. 8 (left panel).
@@
Figure 6: Test trajectory prediction errors for the Mountain Car, Lunar Lander, and Panda environments under unknown dynamics. For these experiments, a Polyflow order of 2 was used for the lifted state construction, and all models were trained with training data subject to 0.1 noise levels. Here, Adam optimizer was replaced by the SOAP optimizer [59] with a learning rate of $3\times 10^{-3}$.
@@ III-D Lunar Lander
The Lunar Lander environment is a Gymnasium environment [53] in which an agent controls the thrusters of a lander to achieve a soft landing on a designated pad. The state space originally comprises the position, velocity, angle, and angular velocity of the lander, while the action space consists of continuous thrust controls.
The test prediction results are presented in fig. 6(b), and the resulting MPC-controlled trajectories are shown in fig. 7(b). The heatmap in fig. 8 (middle panel) further illustrates the model’s robustness under different noise levels and rollout lengths.
@@ III-E Panda Emika Franka Robot
Prediction performance is visualized in fig. 6(c), and the corresponding MPC trajectories are shown in fig. 7(c). Noise sensitivity under varying rollout lengths is depicted in the right panel of fig. 8.
Figure 8: Heatmaps show predictive test performance across different noise levels and sequence lengths. Higher rollout length provides more robustness under noise in training data.
III-F Computational Cost Analysis with Varying Problem Dimensions
To evaluate the computational scalability of our proposed Koopman learning framework, we benchmarked its performance on two nonlinear control environments: Lunar Lander (with a 6-dimensional state and 2-dimensional action space) and Mountain Car (with a 2-dimensional state and 1-dimensional action space).
For each environment, we systematically varied:
  • the Polyflow order ($N_{\phi}$), and
  • the rollout length ($R\in\{1,10,50\}$),
and recorded the training time (in milliseconds) and peak GPU memory usage (in megabytes) per epoch. All experiments were conducted using a single NVIDIA RTX 3080 GPU.
TABLE II: Training Time (ms) / GPU Memory Usage (MB) per Epoch on the Mountain Car Dataset
Polyflow Order R = 1 R = 10 R = 50
2 14 / 26 95 / 102 719 / 443
4 20 / 34 193 / 185 940 / 858
10 25 / 59 337 / 435 1974 / 2106
20 65 / 101 538 / 851 3504 / 4186
TABLE III: Training Time (ms) / GPU Memory Usage (MB) per Epoch on the Lunar Lander Dataset
Polyflow Order R = 1 R = 10 R = 50
2 25 / 27 129 / 112 610 / 491
4 31 / 36 233 / 202 835 / 939
10 36 / 64 496 / 473 2455 / 2288
20 96 / 110 820 / 925 4109 / 4538
As illustrated in Tables II and III, both training time and memory usage increase with higher Polyflow orders and longer rollout horizons. While longer rollouts provide richer temporal context for capturing long-term system dynamics, they also lead to greater computational overhead.
@@ IV Iterative Data Augmentation
Figure fig. 9 visualizes the trajectory of the CartPole system in three different scenarios: no data augmentation, one-time data augmentation, and twice data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. Performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (one-time augmentation) and 67.8 (twice augmentation).
@@
Figure fig. 10 presents violin plots that illustrate the distribution of Model Predictive Control (MPC) costs for the CartPole system on two different tasks, seen and unseen, and iterative models under a noise level of 0.1. The “seen” task refers to scenarios where input bounds are set between [20,20][-20,20], while the “unseen” task refers to tighter input bounds of [10,10][-10,10]. Data augmentation was performed using closed-loop trajectories from the seen task. It is important to note that iterative classical learning [62] cannot be easily generalized to a different task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative data augmentation. The results indicate that both seen and unseen control tasks benefit significantly from iterative data augmentation, with successive iterations leading to lower and more stable MPC costs. This shows the effectiveness of iterative data augmentation in improving the overall performance of the system.
@@ V Ablation Studies
  • Stabilized formulation of Koopman operator in Section section II-G,
@@
V-A Dissipative Koopman Model versus Standard Koopman Model
Figure fig. 11a compares the performance of the dissipative Koopman model and the standard Koopman model in different noise levels for Van der Pol system. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure fig. 12 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the standard Koopman matrix move outside the unit circle, leading to a loss of stability, while the dissipative Koopman model never loses stability during training. Additionally, Figure 13 shows the evolution of the spectral radius ($\rho$, defined as the largest absolute value of the eigenvalues of the Koopman operator) over training steps for the Panda environment, reinforcing the long-term stability of the dissipative Koopman model compared to the standard one.
V-B Learning Rate Scheduler
Figure 11b compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels for Van der Pol system. A cyclic learning rate scheduler varies the learning rate between a predefined minimum and maximum in a periodic manner, often improving convergence stability [63]. For low noise, the cyclic learning rate scheduling achieves a 20% reduction in training error and a 16.67% reduction in test error, with a 9.09% reduction in time taken compared to the constant learning rate. In medium noise scenarios, the cyclic learning rate scheduling shows a 16.67% improvement in training error, 14.29% in test error, and an 8.70% improvement in time taken. However, at high noise levels, cyclic learning rate scheduling underperforms compared to the constant learning rate, with higher training and test errors, as well as slightly longer training times. These results indicate that while the cyclic learning rate scheduler is beneficial at low and medium noise levels, it is less effective in high-noise environments.
V-C Optimizer Switching
Figure fig. 11c examines the effect of switching optimizers on the model’s performance for Van der Pol system. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching from the Adam optimizer to LBFGS after 4000 epochs in a 5000-epoch training cycle resulted in about a 25% error reduction compared to the Adam-only optimizer at low and medium noise levels, demonstrating the efficacy of LBFGS during fine-tuning. However, this advantage diminished at higher noise levels, where the differences between the optimizers became less pronounced. In experiments where optimizers were switched periodically every 500 epochs, results varied, but overall, periodic switching helped minimize the error in certain cases, though it did not consistently outperform the Adam-LBFGS approach. Notably, using only LBFGS led to significantly higher errors, suggesting that while LBFGS is beneficial for fine-tuning, it tends to settle in local minima and does not further improve the error.
V-D Progressive roll-out length versus Constant roll-out length
Figure fig. 11d compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. Progressive roll-out scheduling shows a 20% improvement in mean training errors and a 16.67% improvement in mean test errors compared to the constant roll-out length, indicating better learning and generalization. Although the time taken for training with progressive roll-out scheduling is 6.94% lower, the overall training duration remains slightly higher due to the incremental increase in roll-out length. These results highlight the effectiveness of progressive roll-out scheduling in enhancing model performance.